a chemist needs 100 millimeters of a 76% solution but has only 55% and 85% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
Answer by chessace(471)

Answer x and y in mL
Total: x + y = 100 mL
Mass(whatever): 0.55 x + 0.85 y = 76.0
0.55 x + 0.55 y = 55
.30 y = 21
y = 70 mL of 85%
x = 30 mL of 55%
Which makes sense because 76 is about 2/3 of the way from 55 to 85.